OpenStudy (anonymous):

Find the value of x for this equation 2/x+1 = 4/x-1

OpenStudy (anonymous):

-3

OpenStudy (whpalmer4):

@Ackhat have a look at the code of conduct...you aren't supposed to just give answers \[\frac{2}{x+1} = \frac{4}{x-1}\]You could solve this by cross-multiplying \[2(x-1) = 4(x+1)\]

OpenStudy (anonymous):

thanks